The amount of chlorine evolved by passing 2A of current in an aq. Solution of NACL for 30 minutes is ?
1) 2.64g
2) 1.32g
3) 3.62 g
4) 4.22 g
1
Expert's answer
2019-06-13T10:42:53-0400

Solution.

"m = \frac{M(Cl2) \times I \times t}{n(e) \times F}"

n(e) = 2


"2Cl^- -2e = Cl2^0"

F = 96500

M(Cl2) = 35.5

t = 1800


"m = \frac{35.5 \times 2 \times 1800}{2 \times 96500}"

m = 0.66 g.


"V = \frac{Vm \times I \times t}{n(e) \times F}"

Vm = 22.4 L/mole


"V = \frac{22.4 \times 2 \times 1800}{2 \times 96500}"

V = 0.42 L.

Answer:

m(Cl2) = 0.66 g.

V(Cl2) = 0.42 L.